OpenStudy (anonymous):

Chris eats 3 slices of a pizza that is cut into 8 equal slices. What percentage of the pizza does he eat?

OpenStudy (mertsj):

Don't you think he ate 3/8 of the pizza?

OpenStudy (bahrom7893):

8 => 100% 3 => x%

OpenStudy (bahrom7893):

x = 3*100%/8 = 300/8 %

OpenStudy (bahrom7893):

x=37.5%